OpenStudy (anonymous):

If (goh)(x)=x^2+4x+2 and h(x)=x+1, what is g(x)?

OpenStudy (anonymous):

@terenzreignz , xD etoh pah

OpenStudy (anonymous):

@terenzreignz sawa nah baa?? xDDD

OpenStudy (caozeyuan):

obviously g(x) is of the form ax^2+bx+c, so just subsititude x+1 into that and compare the coefficients after expanding.

OpenStudy (anonymous):

Kindly explain it again, in a way that I can understand it, please

OpenStudy (anonymous):

jeanicca the middle term is 6x

OpenStudy (anonymous):

its not 4x, i think you copied it wrong

terenzreignz (terenzreignz):

Another schoolmate?

OpenStudy (anonymous):

classmate this time. xDD

terenzreignz (terenzreignz):

Fancy.

OpenStudy (caozeyuan):

I'm just saying that g(h(x))=a(x+1)^2+b(x+1)+c, you can find the value for a,b and c by comparing the coeffi.s

OpenStudy (anonymous):

(goh)(x)=x^2+4x+2 and h(x)=x+1 g(x+1)=x^2+4x+2 for x+1 function is x^2+4x+2 replace x by y-1 g(y)=(y-1)^2+4(y-1)+2 =y^2-2y+1+4y-4+2 =y^2+2y-1 replace y by x again g(x)=x^2+2x-1

OpenStudy (anonymous):

Ohhh? Are you sure it's 6x Derek?

OpenStudy (anonymous):

ty jennica

OpenStudy (anonymous):

yup, its 6x

OpenStudy (anonymous):

is the answer ryt have u checked it?

OpenStudy (anonymous):

@ajaysagar thank you so much , you're such a big help :D

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

pls will u try to sort out my question??

OpenStudy (anonymous):

@ajaysagar thank you so much , you're such a big help :D

OpenStudy (anonymous):

ur welcome @Jeannica05

OpenStudy (anonymous):

Derek, I think I'm right regarding dun sa given, baka you're the wrong one kasi ang tinutukoy mo is no. 2 letter c

OpenStudy (anonymous):

ou nga, ahaah, xDD sorry nag rumble utak ko dun, ahahah

OpenStudy (anonymous):

@ajaysagar what sort of question?

OpenStudy (anonymous):

physics oscillations chapter

terenzreignz (terenzreignz):

Labo niyo. :/

OpenStudy (anonymous):

ahah, xD paexplain naman ng maayos ung sagot , xDD di ko maintidihan sagot ehh, xDDD

OpenStudy (anonymous):

hey ty @terenzreignz

OpenStudy (anonymous):

Oh sorry, we haven't tackled that yet and I have no idea about that, sorryy..

terenzreignz (terenzreignz):

Call me TJ @ajaysagar and we'll get along splendidly :)

OpenStudy (anonymous):

kk..

terenzreignz (terenzreignz):

As for you @derek11111 you know that g(h(x)) = x² + 4x + 2 yes?

terenzreignz (terenzreignz):

and that h(x) = x + 1

OpenStudy (anonymous):

yup,

OpenStudy (anonymous):

bakit nga pala g(h(x))?????

terenzreignz (terenzreignz):

definition ng (goh)(x)

OpenStudy (anonymous):

continue po. xD

OpenStudy (anonymous):

@ajaysagar, are you a Filipino?

OpenStudy (anonymous):

kuleet jeanica sinabing ou nga, haahist,m

terenzreignz (terenzreignz):

let's define a new function f(x) = x - 1 You'll notice that this is just the inverse of h(x) = x + 1 (I hope you've mastered finding inverses) So... g(h(f(x))) = goh(f(x)) = ?

terenzreignz (terenzreignz):

So in (goh)(x), replace all x's with x-1

OpenStudy (anonymous):

Di naman sya ung tinanog ko sayo ah si terenzreignz kaya, -_-

OpenStudy (anonymous):

@Jeannica05 i belong to the land of festivals and cultures

terenzreignz (terenzreignz):

Terence @Jeannica05 or TJ. Whichever suits you :P

OpenStudy (anonymous):

ahah, xD nag away nah

OpenStudy (anonymous):

Incredible India

terenzreignz (terenzreignz):

anyway... goh(x) para tanggalin yung h, ang gagawin mo, papalitan mo yung x ng INVERSE NG h. SIGURO NAMAN magaling ka na maghanap ng inverse, kuya? XD @derek11111 :P

OpenStudy (anonymous):

Oww,,,@ajaysagar ok, at derek kasi di ka naman tinatanong eh -_-

terenzreignz (terenzreignz):

Peace guys :)

OpenStudy (anonymous):

hhahaha, xDDD hahahahah nangugulo lang wahahaha

terenzreignz (terenzreignz):

tsk tsk. So ano inverse ng h(x) ? h(x) = x + 1

OpenStudy (anonymous):

Whatever you say derek.

terenzreignz (terenzreignz):

So... everything cool again?

terenzreignz (terenzreignz):

I didn't come here to witness a LQ >:) <grabs popcorn>

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

hala , biglang tumahimik,. xDD

OpenStudy (anonymous):

ganto bah?? xDDD

OpenStudy (anonymous):

\[h ^{-1}(x)=-x+1\]

OpenStudy (anonymous):

Whatever you say derek.

OpenStudy (anonymous):

Oww,,,@ajaysagar ok, at derek kasi di ka naman tinatanong eh -_-

OpenStudy (anonymous):

Di naman sya ung tinanog ko sayo ah si terenzreignz kaya, -_-

OpenStudy (anonymous):

ahah,,xD parang mali tahahaah, xDDDD

terenzreignz (terenzreignz):

mali :P h(x) = x + 1 ganito na lang, ha? isolate x, h(x) - 1 = x tapos yung x yung gawin mong \(h^{-1}\) and yung h(x) gawing x \[x - 1 = h^{-1}(x)\]

OpenStudy (anonymous):

ahhhh, kabaliktaran ng inverse wahahha

terenzreignz (terenzreignz):

kabaliktaran... negative nung inverse. tapos yang inverse na yan (x - 1) yan ang ipasok mo sa goh(x) goh(x) = x² + 4x + 2 makukuha mo yung g(x)

OpenStudy (anonymous):

ahhhh,, \[(x-1)^{2}+4(x-1)+2\]

terenzreignz (terenzreignz):

Yup. simplify tapos yan na ang g(x)

OpenStudy (anonymous):

x^2+4x-1

OpenStudy (anonymous):

x^2+4x+1 atah. xDDD

terenzreignz (terenzreignz):

bawal manghula.

OpenStudy (anonymous):

hahahahah

OpenStudy (anonymous):

g(x)=x^2+4x+1 ayan nah ang ans diba??

terenzreignz (terenzreignz):

mali pa rin :P ano ba ang (x-1)² ?

OpenStudy (anonymous):

x^2+1

terenzreignz (terenzreignz):

<staring in disbelief>

OpenStudy (anonymous):

ai tanga, xDD hahaha

OpenStudy (anonymous):

perfect swuare nga pala

terenzreignz (terenzreignz):

Rectify the situation. S'il vous plaît

OpenStudy (anonymous):

x^2-2x-1

terenzreignz (terenzreignz):

\[\large x^2 -2x \color{red}{+1}\]

OpenStudy (anonymous):

aw, hahhahan sorri typo. xDDDD

terenzreignz (terenzreignz):

right, now simplify this again... \[(x-1)^{2}+4(x-1)+2\]

OpenStudy (anonymous):

x^2-2x+1+4x-4+2

terenzreignz (terenzreignz):

okay... combine like terms.

OpenStudy (anonymous):

x^2-2x+1

OpenStudy (anonymous):

x^2+2x+1

OpenStudy (anonymous):

nakakaduling,waaaa

terenzreignz (terenzreignz):

good... now was that so hard? :P

OpenStudy (anonymous):

ahha, nakakaduling ang maganda kong sulat, ahahahahah

terenzreignz (terenzreignz):

Whatever you say, Derek :D

OpenStudy (anonymous):

ahahahah. xD hala sumbong kita!!! ahhahaha

terenzreignz (terenzreignz):

:3

terenzreignz (terenzreignz):

At least sa akin, may comma :D

OpenStudy (anonymous):

waah!

OpenStudy (anonymous):

xDDDD, tara kain tau, ahahah kakagutom

terenzreignz (terenzreignz):

way ahead of you... nagpadala na ako ... bagal mo kasi sumagot eh :P

OpenStudy (anonymous):

g(x)=x^2+2x-1 ganyan answer ni ajay,,